OpenStudy (anonymous):

find y such that the distance between the points is 8. (0, 0), (3, y) I did SQRT(3^2 + y^2) = 8 y^2 + 9 - 64 = 0 y^2 - 55 = 0

OpenStudy (anonymous):

anyone plz?

OpenStudy (watchmath):

that's correct So \(y=\pm \sqrt{55}\)

OpenStudy (anonymous):

sqrt(3^2 + y^2) = 8 9 + y^2 = 64 y^2 = 55 y = +-sqrt(55)

OpenStudy (anonymous):

allright, thanks